OpenStudy (anonymous):

If y varies inversely as x, find the constant of variation if y = 100 when x = 0.2. A. 100 B. 20 C. 200 D. 10 They all came out no solution @jim

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

y = k/x 100 = k/0.2 k = ??

jimthompson5910 (jim_thompson5910):

you might have been using y = kx, but that only works if y varies directly as x you use y = k/x since y varies inversely as x

OpenStudy (anonymous):

do you dvide

OpenStudy (anonymous):

I got 20

jimthompson5910 (jim_thompson5910):

good k = 100*0.2 k = 20